A life insurance company will sell a PHP 250,000 one-year term life insurance policy for members of armed forces of the Philippines for the premium of PHP 500. Let X denote the net gain from the insurance company based on the collected data of the company, a member of armed police has a 99.96% chance of surviving within one year. Let X the net gain of an insurance company from a single policy holder.


1
Expert's answer
2021-05-19T17:35:05-0400

Survival probability = 0.9996

Death probability = 1-0.9996=0.0004


Probability Distribution of 'X'-


X: C   PHP 250,000

P: 0.9996   0.0004


Expected "E(X)=P_1X_1+P_2X_2"

      "=0.9996C+0.0004(-250000)"

    


For net gain 500, E(X)=500


"\Rightarrow 500 =0.9996C+0.0004(-250000)"


"\Rightarrow C=\dfrac{500+100}{0.9996}=\dfrac{600}{0.9996}=600.24"


Hence Net gain is 600.4